OpenStudy (anonymous):

A flower garden at Dow's Lake is planted in the shape of a parallelogram with a semi-circle at each end. The parallelogram is 5.58 x 2.7 m long and the height of the parallelogram is 2.25 m. What is the area of the garden? If a bag of mulch covers 65 ft2, how many bags of mulch are needed to cover the entire area?

OpenStudy (anonymous):

Area of Parallelogram = height * base = 2.25 * 5.58 = 12.555 m2 Adding two semi circles of opposite sides add to a circle. Hence, there will be two circles of radius 5.58 / 2 = 2.79 m and 2.7 / 2 = 1.35 m. Hence area of these two circles area \[Area = \pi (2.79^{2} + 1.35^{2}) = 30.180 m ^{2}\] Hence total area = 30.180 + 12.555 = 42.735 m2 Now, we know 1 ft = 0.3048 m => \[1 m ^{2} = 10.764 ft ^{2}\] hence, area in ft2 = 459.999 ft2 = 460 ft2 (approx) Hence number of mulch bag required = 460 / 65 = 7.07 = 7 (approx)